OpenStudy (unklerhaukus):

Can it be shown that \[{(\textbf{S~T})}^{-1}= \dots\] \[\dots=(\textbf{T}^{-1}\textbf{S}^{-1})\]

OpenStudy (anonymous):

my head is my only house unless it rains

OpenStudy (unklerhaukus):

* \[{(\textbf{ST})}^{-1}= \dots\] \[\dots=(\textbf{T}^{-1}\textbf{S}^{-1})\]

OpenStudy (anonymous):

it is always true

OpenStudy (anonymous):

\[(ab)^{-1}=b^{-1}a^{-1}\] because \(abb^{-1}a^{-1}=aa^{-1}=I\)

OpenStudy (unklerhaukus):

i dont think i really understand the concept of the inverse of a matrix/ies

OpenStudy (anonymous):

true for functions, elements of a group, anywhere you have inverses. i thas nothing to do with matrices

OpenStudy (unklerhaukus):

why do i have to sow it that way it isnt a logical agrument

OpenStudy (anonymous):

hold on

OpenStudy (unklerhaukus):

it is to do with matrices because the order is what is important

OpenStudy (anonymous):

what is the meaning of \((ab)^{-1}\)? it is the inverse of \((ab)\) i.e. it is the element that when you multiply by \((ab)\) you should get \((ab)(ab)^{-1}=(ab)^{-1}(ab)=I\) the identity

OpenStudy (anonymous):

so no matter what you have, be it numbers, matrices, functions, arbitrary elements of some group, premuations, anything at all it is always true that \((ab)^{-1}=b^{-1}a^{-1}\) always only reason we don't usually see it is because numbers commute whereas functions and matrices do not

OpenStudy (unklerhaukus):

i know it is true i just cant prove it to my self mathematically

OpenStudy (anonymous):

so what is needed to show that \((MN)^{-1}=N^{-1}M^{-1}\) is that when you multiply by \(MN\) you get \(I\)

OpenStudy (anonymous):

and it has to work, because \[MNN^{-1}M^{-1}=MIM^{-1}=MM^{-1}=I\]

OpenStudy (unklerhaukus):

this is true but isn's the same as what i am asking

OpenStudy (unklerhaukus):

why is this the best evidence 'it has to work' because of this other expression is equal to I it just like a convoluted proof that isnt really self justifying

OpenStudy (anonymous):

ok lets ask the question this way. what is the meaning of \((ab)^{-1}\)?

OpenStudy (anonymous):

i am thinking you are looking for some kind of mechanics that will make it work maybe? i am not sure

OpenStudy (anonymous):

suppose we have two functions with inverses say domain all real numbers \(f,g,g^{-1}, f^{-1}\) and composition is defined. then what is the inverse of \(f\circ g\)? i.e. what is \((f\circ g)^{-1}\)?

OpenStudy (unklerhaukus):

To prove similar equation { where the tilde ~ is denoting the transpose \[\widetilde{(\textbf{ST})}=\widetilde{(\textbf{S}_{ik}\textbf{T}_{kj})}_{ij}\]\[= \sum\limits_{k=1}^n s_{jk}t_{ki}\]\[=\sum\limits_{k=1}^n \widetilde{t_{ik}} \widetilde{s_{kj}}\]\[=(\widetilde{ \textbf{T}_{ik}} ~\widetilde {\textbf{S}_{kj} })_{ij}\]\[=(\widetilde{ \textbf{T}}~\widetilde {\textbf{S} })\]

OpenStudy (anonymous):

Well, we are relying on a couple of things for this proof. First, we rely on associativity, that is (ab)(b'a') = a(bb')a' second, we rely on the definition of inverses to get bb' = I, the identity element. That gets us to a*I*a' Then we rely on the definition of the identity element to get a*a' finally we rely on the definition of inverses again to get a*a' = I So we showed that (ab)(b'a') = I Therefore, (b'a') = (ab)'

OpenStudy (unklerhaukus):

i dont know what is \[(f \circ g)^{-1}=\]

OpenStudy (anonymous):

it is \(g^{-1}f^{-1}\)

OpenStudy (anonymous):

nice tilda btw!

OpenStudy (anonymous):

also nice trout mask, i would have used it if i had thought of it first

OpenStudy (unklerhaukus):

Ha ha right right Just dig it That's right "The Mascara Snake" Fast 'n bulbous Tight also

OpenStudy (unklerhaukus):

\[ (\textbf{ST}) {(\textbf{ST})}^{-1} = \textbf {I}\] \[\textbf{T}{(\textbf{ST})}^{-1} = \textbf {S}^{-1}\] \[{(\textbf{ST})}^{-1} = \textbf{T}^{-1}\textbf {S}^{-1}\]